# Two dimensional dimers beyond planarity

### v2

@@ 3.3 From spins to double dimers
The next proposition states a correspondence between observables in our generalised double dimer model and in the spin system which was defined in the Section 2. For each $z\in V$, $u\in\mathbb{N}_{0}$, we define the functions $g_{u,z}:S\mapsto\mathbb{C}$, $g_{z}:S\mapsto\mathbb{C}$, which are
$g_{z}(s)=\begin{cases}\frac{U(0)e^{-2h\cos(rs_{o}^{1})}}{g(s)}&\mbox{ if $z\in V^{e}$}\\ \frac{U(0)e^{-2h\cos(rs_{o}^{1})}}{\overline{g}(s)}&\mbox{ if $z\in V^{o}$}\end{cases},\quad\quad g_{u,z}(s)=\begin{cases}\frac{U(u)}{g(s)}e^{-ui(s^{1}-s^{2})}&\mbox{ if $z\in V^{e}$}\\ \frac{U(u)}{\overline{g}(s)}e^{ui(s^{1}-s^{2})}&\mbox{ if $z\in V^{o}$}\end{cases}.$ (3.8)
The idea is that taking the average of the function $e^{n\,i\,s_{x}^{i}}$, $x\in V$, imposes the presence of $n$ dimers on $\{x,s\}$ (whose colour depends not only on $i\in\{1,2\}$ but also on the parity of $x$ and on the sign of $n$) and may affect the weight which is assigned by the weight function, taking the average of the function $g_{u,x}$ imposes a constraint on the local time at $x$ and the number of $2$-dimers on $\{x,g\}$, while taking the average of the function $g_{x}$ imposes the condition that the local time at $x$ is zero. The domain of the weight function, which has been defined in Section 2, is extended to the negative integers, namely $U:\mathbb{Z}\mapsto\mathbb{R}^{+}_{0}$, under the assumption that $U(n)=0$ for any $n<0$.
Proposition 3.1 (From spins to double dimers).
For $i\in\{1,2\}$, let $u^{i}=(u^{i}_{x})_{x\in V}$, be integer valued vectors and we let $u^{i}_{\pm}=((u^{i}_{\pm})_{x})_{x\in V}$ be such that $(u^{i}_{+})_{x}:=\max\{u^{i}_{x},0\}$ and $(u^{i}_{-})_{x}:=\min\{u^{i}_{x},0\}$. For any $r\in\{1,2\}$, $u\in\mathbb{N}_{0}$, $\beta,h\geq 0$, $A\subset V$, the following identities hold,
@@
$\begin{split}\langle\prod_{x\in V^{e}}e^{i(u^{1}_{x}s_{x}^{1}-u^{2}_{x}s_{x}^{2})}&\prod_{x\in V^{o}}e^{-i(u^{1}_{x}s_{x}^{1}-u^{2}_{x}s_{x}^{2})}\rangle_{G,\beta,h,r}\\ =&\frac{1}{\mathcal{Z}_{G,\beta,h,r}}\mu_{\beta,h,r}\Big{(}\mathbbm{1}_{\{\partial w=(u^{1}_{+}-u^{2}_{-},-u^{1}_{-}+u^{2}_{+})\}}\prod_{x\in V}\frac{U\big{(}k_{x}(w)+(u_{-}^{1})_{x}+(u_{-}^{2})_{x}\big{)}}{U\big{(}k_{x}(w)\big{)}}\Big{)},\end{split}$ (3.9)
$\langle$ $\prod_{z\in A}g_{u,z}(s_{z})\rangle_{G,\beta,h,r}=\mathcal{P}_{G,\beta,h,r}\big{(}\forall z\in A,k_{z}(w)=u\big{)},$ (3.10)
$\langle$ $\prod_{z\in A}g_{z}(s_{z})\rangle_{G,\beta,h,r}=\mathcal{P}_{G,\beta,h,r}\big{(}\forall z\in A,n_{z}(w)=0\big{)}.$ (3.11)
In particular we have that $Z_{G,\beta,h,r}=\mathcal{Z}_{G,\beta,h,r}$.
The proof of the proposition is postponed to Section 8. Note that the weight function in the denominator of (3.9) equals zero only for configurations $w$ such that $\mu_{\beta,h,r}(w)=0$. These do not give any contribution to the expectation, hence (3.9) is well defined.
We say that the function $f:\Omega_{s}\mapsto\mathbb{C}$ is good if it admits a Fourier series with real coefficients and, in addition, $\int_{S^{V}}\boldsymbol{ds}|f(\boldsymbol{s})|<\infty.$ For example, for $A\subset V$, the functions $\prod_{x\in A}\sin(s_{x}^{1})$, $\prod_{x\in A}\cos(s_{x}^{1})$ are good. We say that a function $f:\Omega_{s}\mapsto\mathbb{C}$ is positive good if it is good and, in addition, it admits a Fourier series with non negative coefficients. Applying Proposition 3.1 and our assumptions on $U$ we see that if $f$ is good then $\langle f\rangle_{G,\beta,h,r}$ has a real (finite) value and, if $f$ is positive good, it also satisfies $\langle f\rangle_{G,\beta,h,r}\geq 0.$
@@
We now use the double dimer formulation to show that the average of certain observables for the spin system is zero when $h=0$.
Lemma 3.2.
@@
Suppose that $h=0$. For any $\beta>0$, $x,y\in V$, $n\in\mathbb{Z}\setminus\{0\}$ we have that
$\langle e^{i\,n\,s^{i}_{x}}\rangle_{G,\beta,h,r}$ $=0,$ (3.12)
$\langle e^{ni\,(s^{i}_{x}+s^{i}_{y})}\rangle_{G,\beta,h,r}$ $=0.$ (3.13)
@@
By Proposition 3.1 we observe that $\langle e^{i\,n\,s^{i}_{x}}\rangle$ is a sum over configurations $(m,\pi)\in\Sigma$ such that $\partial m=(n\delta_{x},0)$ or $\partial m=(0,n\delta_{x})$. This condition implies the existence of a path in the extended graph starting with a dimer on the edge $\{x,s\}$ and ending with a dimer on the edge $\{y,g\}$ for some $y\in V$ (indeed, being the graph bipartite, no path starting on $\{x,s\}$ can end with a dimer of the same colour on the same edge). However, any configuration with a strictly positive number of dimers on some edge $\{y,g\}$ gets weight zero, since $h=0$. This concludes the proof of the first identity. We now prove the second identity. Note that by Proposition 3.1 $\langle e^{\pm ni\,(s^{i}_{x}+s^{i}_{y})}\rangle_{G,\beta,h,r}$ is a sum over configurations $(m,\pi)\in\Sigma$ such that if $x$ and $y$ have the same parity then $\partial m=(n\delta_{x}+n\delta_{y},0)$ or $\partial m=(0,n\delta_{y}+n\delta_{x})$, while if $x$ and $y$ have different parity then $\partial m=(n\delta_{x},n\delta_{y})$ or $\partial m=(n\delta_{y},n\delta_{x})$. Being $h=0$, any such configuration gets weight zero. Indeed, being the graph bipartite, no path can end with dimers on the edges $\{x,s\}$ and $\{y,s\}$ if such dimers have the same colour and $x$ and $y$ have the same parity or if such dimers have different colour and $x$ and $y$ have opposite parity. This concludes the proof. ∎
The next proposition relates the two point functions (3.5), which are defined for the generalized double dimer model, to the two-point functions of the spin system appearing in Theorem 2.1.
We say that the weight function $U$ is nice if $U(0)=1$ and $U(n+1)\leq U(n)$ for each $n\in\mathbb{N}_{0}$. We say that the weight function is fast decaying if it is nice and in addition $U(n+1)\leq K\frac{U(n)}{n}$ for each $n\in\mathbb{N}$, where $K_{U}>0$ is an absolute constant. For example, the weight function of the monomer double dimer model is fast decaying, while the weight function of the double dimer model is not even nice.
@@
For any $x,y\in V$ we have that
[…]
where the first inequality holds for any weight function $U$ (satisfying the general assumptions which have been made in Section 2) if $x$ and $y$ have different parity and for any nice weight function $U$ if $x$ and $y$ have the same parity, the second inequality holds only if $U$ is fast decaying.
@@
We omit the subscripts from $\langle\,\,\,\rangle_{G,\beta,0,0}$ for a lighter notation. First note that by Lemma 3.2 we have that
$\langle\cos(\ell s_{x}^{1})\cos(\ell s_{y}^{1})\rangle=\frac{1}{4}\big{(}\langle e^{i\ell(s_{x}^{1}-s_{y}^{1})}\rangle+\langle e^{-i\ell(s_{x}^{1}-s_{y}^{1})}\rangle\big{)}.$
Now if $x$ and $y$ have different parity we obtain from Proposition 3.1 and from the previous expression that
$\langle\cos(\ell s_{x}^{1})\cos(\ell s_{y}^{1})\rangle=\frac{1}{4Z_{G,\beta,0,0}}\Big{(}\sum\limits_{\begin{subarray}{c}(m,\pi)\in\mathcal{W}:\\ \partial m=(\ell\delta_{x}+\ell\delta_{y},0)\end{subarray}}\tilde{\mu}_{\beta,0,0}(m)\prod_{z\in V}\frac{U(n_{z}(m))}{n_{z}(m)!}\\ +\sum\limits_{\begin{subarray}{c}(m,\pi)\in\mathcal{W}:\\ \partial m=(0,\ell\delta_{x}+\ell\delta_{y})\end{subarray}}\tilde{\mu}_{\beta,0,0}(m)\prod_{z\in V}\frac{U(n_{z}(m)-\ell\delta_{x}(z)-\ell\delta_{y}(z))}{n_{z}(m)!}\Big{)}.$ (3.16)
Instead, if $x$ and $y$ have the same parity we obtain from Proposition 3.1 that
$\langle\cos(\ell s_{x}^{1})\cos(\ell s_{y}^{1})\rangle=\frac{1}{4Z_{G,\beta,0,0}}\Big{(}\sum\limits_{\begin{subarray}{c}(m,\pi)\in\mathcal{W}:\\ \partial m=(\ell\delta_{x},\ell\delta_{y})\end{subarray}}\tilde{\mu}_{\beta,0,0}(m)\prod_{z\in V}\frac{U(n_{z}(m)-\ell\delta_{x}(z))}{n_{z}(m)!}\\ +\sum\limits_{\begin{subarray}{c}(m,\pi)\in\mathcal{W}:\\ \partial m=(\ell\delta_{y},\ell\delta_{x})\end{subarray}}\tilde{\mu}_{\beta,0,0}(m)\prod_{z\in V}\frac{U(n_{z}(m)-\ell\delta_{y}(z))}{n_{z}(m)!}\Big{)}.$ (3.17)
If $\ell=1$ we then realise that the first term in the brackets in (3.16) equals $\mu_{G,\beta,0,0}(x\rightarrow y)$. Indeed, any configuration $w=(m,\pi)$ with no dimer touching the ghost vertex and such that $\partial m=(\delta_{x}+\delta_{y},0)$ (with a $1$-dimer on $\{x,s\}$ and on $\{y,s\}$) necessarily has a path from $\{x,s\}$ to $\{y,s\}$ if $x$ and $y$ have different parity. Similarly, using the monotonicity property of $U$, we see that the first term in the brackets in (3.17) is greater or equal than $\mu_{G,\beta,0,0}(x\rightarrow y)$. Indeed, any configuration $w=(m,\pi)$ with no dimer touching the ghost vertex and such that $\partial m=(\delta_{x},\delta_{y})$ (with a $1$-dimer on $\{x,s\}$ and a $2$-dimer on $\{y,s\}$) necessarily has a path from $\{x,s\}$ to $\{y,s\}$ if $x$ and $y$ have the same parity. In both cases, being the second term in the brackets non negative, we conclude the proof of (3.14).
@@
$|f^{-1}(w)|\geq 1,\quad\quad n_{z}(m^{\prime})=n_{z}(m)+\delta_{x}(z)+\delta_{y}(z)\quad\quad\tilde{\mu}_{\beta,h,r}(m)=\mu_{\beta,h,r}(m^{\prime}).$ (3.18)
@@
$\mathcal{G}^{(2)}_{G,\beta}(x,y)=$ $\sum\limits_{w=(m,\pi)\in\{x\leftrightarrow y\}}\tilde{\mu}_{\beta,0,0}(m)\prod_{z\in V}\frac{U(n_{z}(m))}{n_{z}(m)!}$
$=$ $\sum\limits_{w=(m,\pi)\in\{x\leftrightarrow y\}}\frac{1}{|f^{-1}(w)|}\sum\limits_{w^{\prime}=(m^{\prime},\pi^{\prime})\in f^{-1}(w)}\tilde{\mu}_{\beta,0,0}(m)\prod_{z\in V}\frac{U\big{(}n_{z}(m)\big{)}}{n_{z}(m)!}$
$\leq$ $\sum\limits_{w=(m,\pi)\in\{x\leftrightarrow y\}}\sum\limits_{w^{\prime}=(m^{\prime},\pi^{\prime})\in f^{-1}(w)}\tilde{\mu}_{\beta,0,0}(m^{\prime})\prod_{z\in V}\frac{U\big{(}n_{z}(m^{\prime})-\delta_{x}(z)-\delta_{y}(z)\big{)}}{\big{(}n_{z}(m^{\prime})-\delta_{x}(z)-\delta_{y}(z)\big{)}!}$
$\leq$ $K_{U}\,\,\sum\limits_{w^{\prime}=(m^{\prime},\pi^{\prime})\in\mathcal{A}}\tilde{\mu}_{\beta,0,0}(m^{\prime})\prod_{z\in V}\frac{U\big{(}n_{z}(m^{\prime})-2\delta_{x}(z)-2\delta_{y}(z)\big{)}}{n_{z}(m^{\prime})!},$
recognizing that the last term corresponds to the second term in the brackets in (3.16) and then concluding the proof of (3.15).

### v3

@@ 3.3 From spins to double dimers
The next proposition establishes a correspondence between correlations in the multi-occupancy double-dimer model and in the spin system defined in Section 2. In the statement, the domain of the weight function, which has been defined in Section 2, is extended to the negative integers, namely $U:\mathbb{Z}\mapsto\mathbb{R}^{+}_{0}$, under the assumption that $U(n)=0$ for any $n<0$.
Proposition 3.11 (From spin correlations to double dimers).
Let $G$ be a bipartite graph such that $\mathcal{D}_{G}\neq\emptyset$. For $i\in\{1,2\}$, let $u^{i}=(u^{i}_{x})_{x\in V}\in\mathbb{Z}^{V}$ arbitrary vectors, let $u^{i,\pm}=(u^{i,\pm}_{x})_{x\in V}\in\mathbb{Z}^{V}$ be defined as $u^{i,+}_{x}=u^{i}_{x}\vee 0$ and $u^{i,-}_{x}=u^{i}_{x}\wedge 0$ for each $x\in V$. Then, for any $r\in\{1,2\}$, $\beta,h\geq 0$, we have that
$Z^{spin}_{G,\beta,h,r}=Z^{path}_{G,\beta,h,r}>0$
and that
@@
$\Bigl\langle\prod_{x\in V}\prod_{i=1}^{2}(S_{x}^{i})^{u_{x}^{i}}\Bigr\rangle_{G,\beta,h,r}=\frac{1}{{Z}^{path}_{G,\beta,h,r}}\mu_{\beta,h,r}\Bigg(\mathbbm{1}_{\{\partial w=(u^{1,+}-u^{2,-},\,u^{2,+}-u^{1,-})\}}\prod_{x\in V}\frac{U\!\left(k_{x}(w)+u^{1,-}_{x}+u^{2,-}_{x}\right)}{U\!\left(k_{x}(w)\right)}\Bigg),$ (3.11)
The previous proposition shows that averaging the function $(S_{x}^{i})^{n}$, with $n>0$ and $x\in V$, enforces the presence of $n$ $i$-dimers on the edge $\{x,s\}$. Conversely, averaging ${(\overline{S_{x}^{i}})}^{n}=(S_{x}^{i})^{-n}$ enforces the presence of dimers of the opposite colour on $\{x,s\}$, while simultaneously modifying the local weight assigned by the weight function at $x$. The proof of Proposition 3.11 is deferred to Section 8.2. Note that, since the weight of the measure $\mu_{\beta,h,r}$ contain products of factors $U(k_{x}(w))$, the quantity in the RHS of (3.11) is well-defined also when the denominator $U\!\left(k_{x}(w)\right)$ equals zero (division by zero, then, never occurs).
The next proposition states a further correspondence between spin and dimer observables. The proposition involves the following functions, for any $z\in V$,
$\psi_{z}(\boldsymbol{s})=\frac{U(0)\,e^{-2h\cos(rs_{z}^{1})}}{(2\pi)^{2}\gamma_{z}(\boldsymbol{s})}\hskip 18.49988pt\psi_{p,z}(\boldsymbol{s})=\frac{(\overline{S_{z}^{1}S_{z}^{2})}^{p}}{(2\pi)^{2}\gamma_{z}(\boldsymbol{s})}$ (3.12)
with $r,p\in\mathbb{N}_{0}$, whose average induces constraints on the local time.
Proposition 3.12 (Inducing constraints on the local time).
Let $G$ be a bipartite graph such that $\mathcal{D}_{G}\neq\emptyset$. Then, for any $r\in\{1,2\}$, $p\in\mathbb{N}_{0}$, $\beta,h\geq 0$, the following identities hold:
$\Big\langle\prod_{z\in A}\psi_{z}\Big\rangle_{G,\beta,h,r}$ $=P_{G,\beta,h,r}\big(\forall z\in A,\;n_{z}(w)=0\big),$ (3.13)
$\Big\langle\prod_{z\in A}\psi_{p,z}\Big\rangle_{G,\beta,h,r}$ $=P_{G,\beta,h,r}\big(\forall z\in A,\;k_{z}(w)=p\big).$ (3.14)
The proof of the proposition is postponed to Section 8.2 as well.
We now discuss whether the average of a general complex function with respect to our complex spin measure is real and positive.
Definition 3.13 (Real-expectation and Positive-expectation functions).
We say that a function $f:\Omega_{s}\to\mathbb{C}$ is real-expectation if there exists for each $x\in V$ a real-valued sequence $\big(a^{x}_{n_{1},n_{2}}\big)_{n_{1},n_{2}\in\mathbb{Z}}$ such that
$f(s)=\sum_{\boldsymbol{n}\in\mathbb{Z}^{V}\times\mathbb{Z}^{V}}\prod_{x\in V}a^{x}_{n_{x}^{1},n_{x}^{2}}e^{i(n_{x}^{1}s_{x}^{1}-n_{x}^{2}s_{x}^{2})}$
and
$\int_{\Omega_{s}}\boldsymbol{ds}\,|f(\boldsymbol{s})|<\infty.$
We say that $f$ is positive-expectation if it is real-expectation and, in addition, $a^{x}_{n_{1},n_{2}}\geq 0$ for each $n_{1},n_{2}\in\mathbb{Z}$ and $x\in V$.
Applying Proposition 3.11, and recalling that $S_{x}^{i}=e^{\pm is_{x}^{i}}$ we see that, if $f$ is real-expectation, then $\langle f\rangle_{G,\beta,h,r}$ has a real (finite) value and, if $f$ is positive-expectation, it also satisfies $\langle f\rangle_{G,\beta,h,r}\geq 0.$
@@
We now use the double-dimer formulation to show that the average of certain observables for the spin system is zero when the external field is zero.
Proposition 3.14 (Observables with zero average).
@@
Let $G$ be a bipartite graph such that $\mathcal{D}_{G}\neq\emptyset$. Suppose that $h=0$ and let $\beta,r\geq 0$ be arbitrary, fix $n\in\mathbb{N}$ arbitrary. For any $x\in V$ we have that
$\langle({S^{i}_{x}})^{n}\rangle_{G,\beta,h,r}=\langle(\overline{S^{i}_{x}})^{n}\rangle_{G,\beta,h,r}=0.$ (3.15)
Moreover, if $x,y\in V$ have different parity, then
$\langle(\overline{{S^{i}_{x}}}S^{i}_{y})^{n}\rangle_{G,\beta,h,r}=\langle({S^{i}_{x}}\overline{S^{i}_{y}})^{n}\rangle_{G,\beta,h,r}=0,$
while if they have the same parity, then
$\langle({S^{i}_{x}}{S^{i}_{y}})^{n}\rangle_{G,\beta,h,r}=\langle{(\overline{{S^{i}_{x}}{S^{i}_{y}}})}^{n}\rangle_{G,\beta,h,r}=0.$
@@
By Proposition 3.11 we observe that $\langle(\overline{S^{i}_{x}})^{n}\rangle_{G,\beta,h,r}$ and $\langle(S^{i}_{x})^{n}\rangle_{G,\beta,h,r}$ are sums over configurations $w\in\mathcal{W}$ such that $\partial w=(n\delta_{x},0)$ or $\partial w=(0,n\delta_{x})$. Since $h=0$, no path can end at the ghost vertex in such configurations. Any path with both end-steps on $\{x,s\}$ necessarily has dimers of different colour in its last steps. Hence, the corresponding sum is over configurations with zero weight, which proves the first identity.
We now prove the second identity. Since $h=0$, Proposition 3.11 implies that $\langle(\overline{S^{i}_{x}}S^{i}_{y})^{n}\rangle_{G,\beta,h,r}$ and $\langle(S^{i}_{x}\overline{S^{i}_{y}})^{n}\rangle_{G,\beta,h,r}$ are sums over configurations with $n$ dimers on $\{x,s\}$ and $n$ dimers on $\{y,s\}$, where the dimers on $\{x,s\}$ are of different type from those on $\{y,s\}$, and no other dimer touches the source or ghost vertex. This implies the existence of a path whose last steps are dimers of different type on the edges $\{x,s\}$ and $\{y,s\}$. Such a condition can be satisfied only if the distance between $x$ and $y$ is even; otherwise the sum is empty. This proves the second identity.
Finally, we prove the third identity. If $x=y$ then the first identity applies, so we assume $x\neq y$. Again with $h=0$, Proposition 3.11 implies that $\langle(\overline{S^{i}_{x}S^{i}_{y}})^{n}\rangle_{G,\beta,h,r}$ and $\langle({S^{i}_{x}S^{i}_{y}})^{n}\rangle_{G,\beta,h,r}$ are sums over configurations with $n$ dimers on $\{x,s\}$ and $n$ dimers on $\{y,s\}$, all of the same type, and no further dimer touching the source or ghost vertex. This implies the existence of a path whose last steps are dimers of the same type on the edges $\{x,s\}$ and $\{y,s\}$. Such a condition can be satisfied only if the distance between $x$ and $y$ is odd; otherwise the sum is empty. This proves the third identity.
The next proposition relates the two point functions (3.6), which are defined for the multi-occupancy double-dimer model, to the two-point functions of the spin system appearing in Theorem 2.2.
We say that the weight function $U$ is normalized monotone if $U(0)=1$ and $U(n+1)\leq U(n)$ for each $n\in\mathbb{N}_{0}$. We say that the weight function is fast decaying if it is normalized monotone and, in addition, $U(n+1)\leq K_{U}\frac{U(n)}{n}$ for each $n\in\mathbb{N}$, where $K_{U}>0$ is an absolute constant. For example, the weight function of the monomer double-dimer model is fast decaying, while the weight function of the double-dimer model is not even normalized monotone, since in this case $U(0)=0$.
@@
Let $G$ be a bipartite graph such that $\mathcal{D}_{G}\neq\emptyset$. For any $x,y\in V$ we have that
$\mbox{ if $x\neq y$}\qquad\qquad\mathcal{G}_{G,\beta}^{(1)}(x,y)$ $\leq 4\,\langle\cos(s_{x}^{1})\,\cos(s_{y}^{1})\rangle_{G,\beta,0,0},$ (3.16)
$\mbox{ if $d(x,y)\in 2\mathbb{N}+1$}\qquad\qquad\mathcal{G}_{G,\beta}^{(2)}(x,y)$ $\leq K_{U}\,\langle\cos(2s_{x}^{1})\,\cos(2s_{y}^{1})\rangle_{G,\beta,0,0},$ (3.17)
where the first inequality holds for any weight function $U$ (satisfying the general assumptions which have been made in Section 2) if $x$ and $y$ have different parity and for any normalized monotone weight function $U$ if $x$ and $y$ have the same parity, the second inequality holds only if $U$ is fast decaying.
@@
We omit the subscripts from $\langle\,\,\,\rangle_{G,\beta,0,0}$ for a lighter notation. First note that by the definition of the spin variables (2.1) we have
$\langle\cos(\ell s_{x}^{1})\cos(\ell s_{y}^{1})\rangle=\frac{1}{4}\,\big\langle\big((S_{x}^{1})^{\ell}+(\overline{S_{x}^{1}})^{\ell}\big)\big((S_{y}^{1})^{\ell}+(\overline{S_{y}^{1}})^{\ell}\big)\big\rangle.$
Now if $x$ and $y$ have different parity then by Proposition 3.14 we deduce that $\langle\cos(\ell s_{x}^{1})\cos(\ell s_{y}^{1})\rangle=\frac{1}{4}{\langle(S_{x}^{1}S_{y}^{1})}^{\ell}\rangle+\frac{1}{4}\langle{(\overline{S_{x}^{1}S_{y}^{1}})}^{\ell}\rangle$. We then obtain from Proposition 3.11 that
$\langle\cos(\ell s_{x}^{1})\cos(\ell s_{y}^{1})\rangle=\frac{1}{4Z^{path}_{G,\beta,0,0}}\Big(\sum\limits_{\begin{subarray}{c}w\in\mathcal{W}:\\ \partial w=(\ell\delta_{x}+\ell\delta_{y},0)\end{subarray}}\nu_{\beta,0,0}(m)\prod_{z\in V}\frac{U(n_{z}(m))}{n_{z}(m)!}\\ +\sum\limits_{\begin{subarray}{c}w\in\mathcal{W}:\\ \partial w=(0,\ell\delta_{x}+\ell\delta_{y})\end{subarray}}\nu_{\beta,0,0}(m)\prod_{z\in V}\frac{U(n_{z}(m)-\ell\delta_{x}(z)-\ell\delta_{y}(z))}{n_{z}(m)!}\Big).$ (3.18)
Instead, if $x$ and $y$ have the same parity, we obtain from Proposition 3.14 that $\langle\cos(\ell s_{x}^{1})\cos(\ell s_{y}^{1})\rangle=\frac{1}{4}\langle{(S_{x}^{1}\overline{S_{y}^{1}})}^{\ell}\rangle+\frac{1}{4}\langle{(\overline{S_{x}^{1}}S_{y}^{1})}^{\ell}\rangle$ and from Proposition 3.11 that
$\langle\cos(\ell s_{x}^{1})\cos(\ell s_{y}^{1})\rangle=\frac{1}{4Z^{path}_{G,\beta,0,0}}\Big(\sum\limits_{\begin{subarray}{c}(m,\pi)\in\mathcal{W}:\\ \partial m=(\ell\delta_{x},\ell\delta_{y})\end{subarray}}\nu_{\beta,0,0}(m)\prod_{z\in V}\frac{U(n_{z}(m)-\ell\delta_{x}(z))}{n_{z}(m)!}\\ +\sum\limits_{\begin{subarray}{c}(m,\pi)\in\mathcal{W}:\\ \partial m=(\ell\delta_{y},\ell\delta_{x})\end{subarray}}\nu_{\beta,0,0}(m)\prod_{z\in V}\frac{U(n_{z}(m)-\ell\delta_{y}(z))}{n_{z}(m)!}\Big).$ (3.19)
If $\ell=1$ we then realise that the first term in the brackets in (3.18) equals $\mu_{\beta,0,0}(x\rightarrow y)$. Indeed, any configuration $w=(m,\pi)$ with no dimer touching the ghost vertex and such that $\partial m=(\delta_{x}+\delta_{y},0)$ (with a $1$-dimer on $\{x,s\}$ and on $\{y,s\}$) necessarily has a path from $\{x,s\}$ to $\{y,s\}$ if $x$ and $y$ have different parity. Similarly, using the monotonicity property of $U$, we see that the first term in the brackets in (3.19) is greater or equal than $\mu_{\beta,0,0}(x\rightarrow y)$. Indeed, any configuration $w=(m,\pi)$ with no dimer touching the ghost vertex and such that $\partial m=(\delta_{x},\delta_{y})$ (with a $1$-dimer on $\{x,s\}$ and a $2$-dimer on $\{y,s\}$) necessarily has a path from $\{x,s\}$ to $\{y,s\}$ if $x$ and $y$ have the same parity. In both cases, being the second term in the brackets non negative, we conclude the proof of (3.16).
@@
$|f^{-1}(w)|\geq 1,\qquad\qquad n_{z}(m^{\prime})=n_{z}(m)+\delta_{x}(z)+\delta_{y}(z)\qquad\qquad\nu_{\beta,h,r}(m)=\nu_{\beta,h,r}(m^{\prime}).$ (3.20)
@@
$\mathcal{G}^{(2)}_{G,\beta}(x,y)=$ $\sum\limits_{w=(m,\pi)\in\{x\leftrightarrow y\}}\nu_{\beta,0,0}(m)\prod_{z\in V}\frac{U(n_{z}(m))}{n_{z}(m)!}$
$=$ $\sum\limits_{w=(m,\pi)\in\{x\leftrightarrow y\}}\frac{1}{|f^{-1}(w)|}\sum\limits_{w^{\prime}=(m^{\prime},\pi^{\prime})\in f^{-1}(w)}\nu_{\beta,0,0}(m)\prod_{z\in V}\frac{U\big(n_{z}(m)\big)}{n_{z}(m)!}$
$\leq$ $\sum\limits_{w=(m,\pi)\in\{x\leftrightarrow y\}}\sum\limits_{w^{\prime}=(m^{\prime},\pi^{\prime})\in f^{-1}(w)}\nu_{\beta,0,0}(m^{\prime})\prod_{z\in V}\frac{U\big(n_{z}(m^{\prime})-\delta_{x}(z)-\delta_{y}(z)\big)}{\big(n_{z}(m^{\prime})-\delta_{x}(z)-\delta_{y}(z)\big)!}$
$\leq$ $K_{U}\,\,\sum\limits_{w^{\prime}=(m^{\prime},\pi^{\prime})\in\mathcal{A}}\nu_{\beta,0,0}(m^{\prime})\prod_{z\in V}\frac{U\big(n_{z}(m^{\prime})-2\delta_{x}(z)-2\delta_{y}(z)\big)}{n_{z}(m^{\prime})!},$
where $K_{U}>0$ denotes the decay constant associated with the fast–decaying weight function $U$. Observing that the last term coincides with the second term inside the brackets in (3.18) when $\ell=2$, we thereby complete the proof of (3.17). ∎